Find the second and third derivatives of the following functions by using Matlab

1. f(x) = 3cosx-2sinx x² 2. f(x) = ln(x) - 1 3. f(x) = 5x3 - 2x2 + 7x + 10
Answer #1

eig.mx mdof.mx partCQ2.mx 1 4 clc syms x f = (3* cos(x)-2*sin(x))/(x^2); % second derivative d2 = diff(f,2) % third derivativ

+ eig.mx mdot.mx partCQ2.mx - 3 1 4 1 clc 2 syms x f = log(x)-1; % second derivative d2 = diff(f,2) % third derivative 7 - d3

1 clc 2 syms x 3 f = 5*x^3-2*x^2+7*x+10; 4 % second derivative 5 d2 = diff(f,2) % third derivative 7 d3 = diff(f,3) 8 COMMAND

Feel free to ask for further clarification.

please up-vote if you liked this solution. As it will encourage me to solve more and more problems and help students.

Thank you,